how do you write this in vertex form??
y= x{squared}+ 18x-4

it is algebra 2

Answer by nerdybill(7384)

The vertex form is:
y= a(x-h)^2+k
where
(h,k) is the vertex
.
For more info, see:
http://www.mathwarehouse.com/geometry/parabola/standard-and-vertex-form.php
.
For your problem:
y= x^2 + 18x - 4
.
You must "complete the square":
y= (x^2 + 18x) - 4
Take half of 18 and square it:
y= (x^2 + 18x + 81) - 4 - 81
y= (x^2 + 18x + 81) - 85
y = (x+9)^2 - 85 (this is what they're looking for)